OpenStudy (anonymous):

the perimeter of a basketball court is 102 meters and the length is 6 meters longer than twice the width. what are the length and width?

OpenStudy (anonymous):

which one are you in the pic

OpenStudy (anonymous):

the girl. lol

OpenStudy (anonymous):

perimeter is length+length+width+width the length is 6, twice the width. so 6=2 X ?

OpenStudy (anonymous):

hmm so 2L X 2W=102?

OpenStudy (anonymous):

Perimiter= 2(width) + 2(length) 102=2W+2L The length is 6 meters more than twice the width L=2W+6 Now put this value in for L So... 2(2W+6) + 2W= 102 4W+12+2W=102 6W+12=102 6W=90 w=15 Now we know the width... Length= 2W+6 (from before) L= 2(15)+6 L=30+6 L=36 Check.... 2L+2W=102 2(36)+2(15)=102 :)

OpenStudy (anonymous):

length=6+2W perimeter= 2L + 2W so substitute our length formula in for L and 102 in for perimeter like so 102=2(6+2W) + 2W which would give you?

OpenStudy (anonymous):

i dont know why im still confused so 2 x length and 2 x width =102. and then u add 6 to width?

OpenStudy (anonymous):

Yes, we want to get the equation into a form that only has one variable. In this case, they give us the hint. When they say "The length is 6 meters longer than twice the width" We can write like this.. Length=6 meters longer then twice the width L=6+2W Now we have the length in terms of the width and we can rewrite the equation all in terms of the width and therefore solve for it. Once we know the width, we use the same piece of info to figure out the length (just plug the value of W and solve for L) Hope that helps!!!